OpenStudy (anonymous):

Can anyone help me with subtle U-substitution?? integral of e^(2z)/e^(2z)-4e^(-2z) dz

OpenStudy (anonymous):

\[\int\limits \frac{ e ^{2z} }{ e^{2z}-4e^{-2z} } dz\]

random231 (random231):

take e^2z as x

OpenStudy (anonymous):

as x?

random231 (random231):

yeah watever you are comfortable with

zepdrix (zepdrix):

ya i like that idea. Confused on how it will work out rod?

hartnn (hartnn):

would the substitution \(\Large e^{4z} = x\) after multiplying numerator and denominator by \(e^{2x}\) make the solution easier?

OpenStudy (anonymous):

Yes, I'm not sure where I would go from there

hartnn (hartnn):

i meant ....by \(e^{2z}\)

zepdrix (zepdrix):

\[\large\rm x=e^{2z}\qquad\qquad\qquad\frac{1}{x}=e^{-2z}\]Yah multiplying through is another option :)

random231 (random231):

why not? exactly zepdrix. youll get integral 1/(x^2+2^2) form

hartnn (hartnn):

with the substitution \(x= e^{2z}\), you will get a quadratic term in the denominator, with the substitution \(x= e^{4z}\), you will get a linear term in the denominator,

random231 (random231):

but then what about the numerator? it will be like integral x^2/(x+4)?

zepdrix (zepdrix):

You figure this one out ruby rod? :)

hartnn (hartnn):

@random231 the numerator will be \(e^{4z} dz\) which will actually be dx so your new integral will be of the form \(\dfrac{dx}{x-4}\)

random231 (random231):

mhmm yeah hartnn that wub be a lot easier!

OpenStudy (anonymous):

I think I'm starting to understand it, I'm only wondering what it was multiplied by to simplify the denominator that way?

hartnn (hartnn):

denominator: \(\Large e^{2z}-4e^{-2z} = e^{2z}-4{\dfrac{1}{e^{2z}}} = \dfrac{e^{4z}-4}{e^{2z}}\)

OpenStudy (anonymous):

ohhh ok, now I see it

hartnn (hartnn):